OpenStudy (anonymous):

What is the volume of the trapezoidal right prism? A. 846 m3 B. 720 m3 C. 423 m3 D. 56 m3

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

420?

OpenStudy (igreen):

I don't think so.

OpenStudy (igreen):

\(V = \dfrac{1}{2}h(a + b)\) Where 'h' is the height and 'a' and 'b' are the two base lengths.

OpenStudy (anonymous):

Blonde moment! xD Lol

OpenStudy (igreen):

\(V = \dfrac{1}{2}(6)(7 + 40)\)

OpenStudy (anonymous):

I just did that and I got 141, thats not even an option. :/

OpenStudy (igreen):

Hmm..

OpenStudy (igreen):

Oh..then we have to multiply it by the length..3.

OpenStudy (igreen):

141 * 3 = ?

OpenStudy (igreen):

\(V = \dfrac{1}{2}h(a + b) \times l\) Where 'h' is the height, 'a' and 'b' are the two base lengths, and 'l' is the length.

OpenStudy (anonymous):

Yes, thats option C.

OpenStudy (destanie_98):

c

OpenStudy (igreen):

Yep, you got it \(\Huge\ddot\smile\)

OpenStudy (anonymous):

Thanks!! :)

OpenStudy (anonymous):

O.O I did this and came to see if i was correct and 100% Thanks!